find the vapor pressure of a solution containing 35.0g calcium chloride (CaCl2) in 750.0g water. Note that the P water= 47.1 torr at 37.0C


1
Expert's answer
2022-03-07T17:26:03-0500

"PV= nRT"

Volume = 3.5× 10^-5

Moles= mass/ rfm= 35/111= 0.315moles

T= 310.15K

P× 3.5 × 10^-5= 0.315× 8.314× 310.15

P= 23196060 torr